# RooseBERT: A New Deal For Political Language Modelling

### v1

@@ 4 Experimental Setup
Each PLM was fine-tuned on the four downstream tasks described in Section 3. During fine-tuning, a sequence classification head was added on top of each model for the Sentiment Analysis and Relation Classification tasks, while a token classification head was used for the NER and Argument Detection tasks. Following the approach of Devlin et al. (2019), we experimented with various hyper-parameters: learning rates of $\{2,3,5\}\times e^{-5}$; weight decay values of $0.1$ and $0.01$; batch sizes of $8$, $16$, and $32$; number of epochs of $2$, 33, or $4$; and maximum sequence lengths of $128$, $256$, and $512$ tokens. Due to constraints arising from the computational power required to train the large architectures, for the fine-tuning of RooseBERT-Large, we fixed the max sequence length to $256$ tokens and the weight decay to $0.01$ while leaving other hyper-parameters unaltered. For all tasks, we used the original train, development and testing splits provided with each dataset by the authors.
@@
The performances on the downstream tasks of RooseBERT-cont and RooseBERT-scr (base architecture) were compared to the performances of BERT-base (used as a baseline) and of similar political PLMs, ConfliBERT and PoliBERTweet, each fine-tuned on the downstream tasks with the same hyper-parameters as RooseBERT. However, since PoliBERTweet was pre-trained on political tweets with a maximum sequence length of 128, we fixed the sequence length to 128 for this model during fine-tuning.
To the best of our knowledge, no politics-based language models trained with a large architecture is available in literature for comparison with RooseBERT-large. Therefore, we used BERT-Large as the sole comparison. For each run, we used one A100 GPU and the Transformer Library of HuggingFace with Python 3.10.
@@ 5 Evaluation
Pre-Training
Perplexity (PPL) (Jelinek et al. 2005) was used as a metric to determine the best performing pre-trained model given the hyper-parameters previously described. Perplexity is defined as the exponentiated average negative log-likelihood of a sequence $X~=~(x_{0},x_{1},x_{2}..x_{t})$ as follows: $PPL(X)=exp\{-\frac{1}{t}\sum^{t}_{i}\ log\ p_{\theta}(x_{i}|x_{<i})\}$. Perplexity was used to evaluate the model’s ability to predict masked tokens on an unseen test dataset representative of the overall corpus distribution. A lower perplexity indicates that the model assigns higher probabilities to the correct next words, i.e., it is more confident in its predictions, less confused by the text.
Base Architecture.
Table 2 reports the perplexity scores achieved by each model, including RooseBERT, baselines, and related models. For conciseness, we only report the best-performing configurations for RooseBERT. Specifically, for both RooseBERT-cont and RooseBERT-scr, a learning rate of $1e^{-4}$ consistently yielded the lowest perplexity on an held out portion of the testing dataset (see Section 3). Final evaluation loss also confirmed this trend. For the cased model, the lowest loss was 0.899 with $1e^{-4}$, followed by 0.901 with $3e^{-4}$, while all other rates (i.e., $5e^{-4}$, $\{1,3,5\}\times e^{-5}$) resulted in losses exceeding 0.925. For the uncased model, the lowest loss was 0.924 with $1e^{-4}$, increasing to 0.933 with $3e^{-4}$, and surpassing 0.951 for the remaining learning rates.
@@
In terms of training time, training RooseBERT-cont-cased and RooseBERT-cont-uncased model required 18 hours for 150k steps. Training the scr model as considerably more costly, with a total training time of 31 hours and 250k steps.
Large Architecture.
We evaluated the perplexity of three large architectures (BERT-large, RooseBERT-large-cont, RooseBERT-large-scr) on a held-out portion of the testing dataset. The BERT-large model, lacking domain-specific pre-training, performed poorly, with perplexity values even higher than those of its base counterpart: 103,121.090 (cased) and 37,326.050 (uncased). In contrast, both RooseBERT variants significantly outperformed BERT, with RooseBERT-large-cont achieving the lowest perplexity: 2.274 (cased) and 2.355 (uncased). RooseBERT-large-scr achieved 2.601 in its cased version and 2.643 in its uncased version. Results confirm that continued pre-training (cont) is more effective than from-scratch training (scr) in this context, allowing the model to retain general linguistic knowledge while adapting to domain-specific language.
The hyper-parameters used to train the large version of RooseBERT are the same used to train the base version of RooseBERT with the exception of the number of steps that in the large version is of 125K. Training the large architecture required more computational power with respect to the base architecture, with a total training time of 35 hours for both the cont and scr versions. As a consequence, due to time constraints, it was only possible to train the large versions of RooseBERT for 125K steps.
Fine-Tuning
Base Architecture.
BERT, RooseBERT, ConfliBERT and PoliBERTweet available versions (cont, scr, cased and uncased) were fine-tuned on four downstream tasks using the hyper-parameters declared in Section 4. In particular, ConfliBERT is available in both cased and uncased version, with cont and scr configurations. PoliBERTweet is only available in cased version with the cont configuration. The results of the fine-tuning of the base architectures are reported in Table 4. Each reported score represents the average performance over 5 independent runs, each initialised with different seeds.
@@
F1 Score $(\uparrow)$
@@
Macro F1 Score $(\uparrow)$
BERT-base-cased 0.528 $\pm$ 0.010 0.655 $\pm$ 0.044 0.464 $\pm$ 0.003 0.638 $\pm$ 0.025
BERT-base-uncased 0.504 $\pm$ 0.009 0.690 $\pm$ 0.003 0.466 $\pm$ 0.003 0.662 $\pm$ 0.007
ConfliBERT-cont-cased 0.441 $\pm$ 0.014 0.683 $\pm$ 0.011 0.460 $\pm$ 0.002 0.659 $\pm$ 0.006
ConfliBERT-cont-uncased 0.492 $\pm$ 0.012 0.691 $\pm$ 0.000 0.464 $\pm$ 0.003 0.659 $\pm$ 0.007
ConfliBERT-scr-cased 0.518 $\pm$ 0.004 0.675 $\pm$ 0.012 0.459 $\pm$ 0.005 0.621 $\pm$ 0.051
ConfliBERT-scr-uncased 0.516 $\pm$ 0.015 0.693 $\pm$ 0.006 0.458 $\pm$ 0.002 0.659 $\pm$ 0.007
PoliBERTweet 0.338 $\pm$ 0.039 0.654 $\pm$ 0.024 0.443 $\pm$ 0.004 0.667 $\pm$ 0.004
RooseBERT-cont-cased 0.477 $\pm$ 0.020 0.726 $\pm$ 0.009 0.473 $\pm$ 0.001 0.663 $\pm$ 0.007
RooseBERT-cont-uncased 0.389 $\pm$ 0.072 0.726 $\pm$ 0.005 0.466 $\pm$ 0.006 0.611 $\pm$ 0.109
RooseBERT-scr-cased 0.499 $\pm$ 0.007 0.725 $\pm$ 0.007 0.477 $\pm$ 0.005 0.669 $\pm$ 0.011
RooseBERT-scr-uncased 0.484 $\pm$ 0.007 0.728 $\pm$ 0.004 0.480 $\pm$ 0.001 0.674 $\pm$ 0.003
Table 4: Average and standard deviation results of the downstream tasks for the base architectures.
RooseBERT, especially in its scr version, outperforms its competitors across all tasks, with the exception of NER. This exception can be attributed to the nature of the CrossNER dataset used for the NER evaluation. Unlike the other datasets, which are tailored to political debates, CrossNER focuses more broadly on the political domain. As a result, it may not align as closely with RooseBERT’s debate-specific training (see Section 3).
In sentiment analysis, RooseBERT-scr-uncased achieves a 3% improvement over best-performing baseline and competitors. This gain is statistically significant, as demonstrated by paired t-tests conducted against BERT-base-uncased (p=0.00003) and ConfliBERT-scr-uncased (p=0.00062). In Argument Component Detection, RooseBERT-scr-uncased achieves a 2% improvement over best-performing baselines (BERT-base-uncased, p=0.00022) and competitors (ConfliBERT-cont-uncased, p=0.00034). In Argument Relation Prediction, RooseBERT-scr-uncased improves over of 1% over best performing baselines (BERT-base-uncased, p=0.00664) and competitors (PoliBERTweet, p=0.01922). These results emphasise the importance of domain-specific pre-training, where task-relevant embeddings contribute to a more refined understanding of argumentative structure.
Large Architecture.
To the best of our knowledge, the current state of the art does not include pre-trained language models for political debates which utilise a large architecture. Therefore, the only possible comparison for RooseBERT-large is its baseline, BERT-large. Table 5 reports the results of the fine-tuning of BERT-large and RooseBERT-large on the four downstream tasks and hyper-parameters discussed in Section 4. Each result is averaged over five runs with different random seeds.
Models
NER
Macro F1 Score $(\uparrow)$
Sentiment Analysis
F1 Score $(\uparrow)$
Arg. Component Detection
Macro F1 Score $(\uparrow)$
Arg. Relation Prediction
Macro F1 Score $(\uparrow)$
BERT-large-cased 0.594 $\pm$ 0.021 0.691 $\pm$ 0.000 0.469 $\pm$ 0.005 0.576 $\pm$ 0.198
BERT-large-uncased 0.492 $\pm$ 0.152 0.680 $\pm$ 0.025 0.466 $\pm$ 0.002 0.410 $\pm$ 0.283
RooseBERT-large-cont-cased 0.553 $\pm$ 0.006 0.718 $\pm$ 0.008 0.475 $\pm$ 0.002 0.499 $\pm$ 0.264
RooseBERT-large-cont-uncased 0.419 $\pm$ 0.214 0.715 $\pm$ 0.013 0.410 $\pm$ 0.135 0.598 $\pm$ 0.210
RooseBERT-large-scr-cased 0.572 $\pm$ 0.010 0.717 $\pm$ 0.004 0.475 $\pm$ 0.002 0.687 $\pm$ 0.008
RooseBERT-large-scr-uncased 0.523 $\pm$ 0.009 0.718 $\pm$ 0.003 0.479 $\pm$ 0.002 0.695 $\pm$ 0.006
In this case, we again observe the superiority of the scr configuration over the cont one, with RooseBERT-large-scr-uncased outperforming all other models on three out of four tasks. Notably, in the large architecture, the cont configuration appeared more prone to overfitting under certain random seeds, especially in tasks such as argument detection and relation classification, leading to inconsistent or degraded performance.
Performance gains from scaling up to the large architecture were modest when compared to the base model. A notable exception is relation classification, where RooseBERT-large achieved a Macro F1 score of 0.695, improving over the base model’s score of 0.674. This findings suggests that, in some cases, a well-optimised base model is sufficient to capture most of the relevant patterns in the data, and that increasing model size does not always guarantee substantial improvements—particularly when data complexity or volume does not fully leverage the larger parameter space.
Ablation Study
@@
To better understand which portions of the dataset contribute most to our results, we conducted an ablation study. However, since the dataset consists of eight subsets, performing a leave-one-out ablation would have been computationally expensive. Instead, we clustered the dataset subsets based on the similarity of their content. This was achieved using cosine similarity computed over the average of the embeddings of the sentences of each dataset. To obtain the sentence embeddings we used all-mpnet-base-v2222huggingface.co/sentence-transformers/all-mpnet-base-v2 Sentence Transformer. Figure 1 shows the results of the clustering, performed using Agglomerative Clustering with a distance threshold of 0.2.
Figure 1: Cluster datasets.
Models
NER
Macro F1 Score $(\uparrow)$
Sentiment Analysis
F1 Score $(\uparrow)$
Arg. Component Detection
Macro F1 Score $(\uparrow)$
Arg. Relation Prediction
Macro F1 Score $(\uparrow)$
Cluster#0 0.491 $\pm$ 0.015 0.733 $\pm$ 0.009 0.475 $\pm$ 0.003 0.667 $\pm$ 0.005
Cluster#1 0.484 $\pm$ 0.014 0.732 $\pm$ 0.009 0.477 $\pm$ 0.003 0.676 $\pm$ 0.005
Cluster#2 0.492 $\pm$ 0.008 0.725 $\pm$ 0.004 0.477 $\pm$ 0.002 0.669 $\pm$ 0.009
Cluster#3 0.465 $\pm$ 0.014 0.734 $\pm$ 0.002 0.477 $\pm$ 0.002 0.663 $\pm$ 0.004
RooseBERT-scr-uncased 0.484 $\pm$ 0.007 0.728 $\pm$ 0.004 0.480 $\pm$ 0.001 0.674 $\pm$ 0.003
@@
As shown in Figure 1, four distinct clusters emerge. One cluster contains only U.S. televised debates (cluster#0), which is expected since they are the only ones not held in a parliamentary setting. Another cluster groups the General Debate of the United Nations (UNGDC) with the United Nations Security Council debates (cluster#1), reflecting their shared institutional origin and likely overlap in terminology and thematic content. The remaining two clusters consist, respectively, of UK Parliament debates on one side (cluster#2), and debates from the European, Australian, Scottish, and more recent UK parliaments on the other (cluster#3). This separation may stem from the broader temporal span of the UK Parliament data (1979–2019), while the other subsets mostly cover the period from 1998 onward. Supporting this interpretation, the latter cluster also includes a subset of UK debates from 2009–2019, which are more aligned with the time frame and content of the other parliamentary sources. Once the clusters were identified, we pre-trained four separate RooseBERT models, each using a leave-one-out strategy with respect to the clusters. For instance, the model labelled Cluster#1 was trained on all debates except those in Cluster 1 (i.e., UNGDC and UN Security Council debates). This setup allows us to assess the individual contribution of each cluster to the overall performance by observing the impact of its exclusion.
@@
As shown in Table 7, all models (except the one trained without the US Debates) perform worse than the original model, confirming that most subsets contribute positively to the model’s overall knowledge. The exclusion of US Debates appears to have little impact on perplexity, likely because these debates differ in nature from the rest of the dataset, which is predominantly composed of parliamentary debates.
@@
Table 6 shows the results of the fine-tuning of these models on the four different downstream tasks. Results are averaged on a total of 5 runs with different seeds. The performance of the models trained with different dataset configurations remains largely consistent across all downstream tasks, showing minimal to no degradation compared to the original RooseBERT model. This consistency highlights RooseBERT’s strong generalisation capabilities. Notably, we hypothesised a performance drop in the argument component detection and relation prediction tasks for the model trained without Cluster#0, as these tasks are based on U.S. political debates, i.e., the portion excluded from that model’s pre-training. However, the results show no significant decrease, suggesting that the structural patterns required for these tasks can be effectively learned from other parliamentary debates. This finding indicates that the model can generalise the discourse structure of political debates even when a key domain is withheld during pre-training.
Comparison with Large Language Models
To compare RooseBERT with Large Language Models (LLMs), we conducted two evaluations: (1) fine-tuning with a token or sequence classification head, mirroring RooseBERT’s setup; and (2) prompt-based zero and few-shot (3-shot) inference using natural language prompts without architectural changes. We evaluated gemma-3-4b-it, Mistral-7B-Instruct-v0.3, Llama-3.1-8B-Instruct, and gpt-4.1-nano, using 0.0 as temperature for determinism. LLMs severely underperformed in the fine-tuning setup, showing some improvement in prompt-based settings. RooseBERT consistently outperformed all models, particularly in the argument component detection and classification task, Gemma achieved a Macro F1 of 0.219 (0-shot). In the sentiment analysis task, Llama performed best with 0.597 F1 (3-shot); in the NER task, Llama performed best with 0.390 Macro F1 (3-shot); in the argument relation prediction and classification task, Mistral performed best with 0.446 Macro F1 (0-shot). See supplementary material for full results.

### v2

@@ 4.2 Fine-tuning Setup
Each PLM was fine-tuned on the four downstream tasks described in Section 3. Following the approach of Devlin et al. (2019), we experimented with various hyper-parameters: learning rates of $\{2,3,5\}\times e^{-5}$; batch sizes of $8$, 1616, and $32$ and number of epochs of $2$, 33, or $4$. For each model, we used its maximum available sequence length. For all tasks, we used the original train, development and testing splits provided with each dataset.
@@
The performances on the downstream tasks of RooseBERT were compared to the performances of BERT-base (used as a baseline) and of similar political PLMs, ConfliBERT and PoliBERTweet, each fine-tuned on the downstream tasks with the same hyper-parameters as RooseBERT.
For each run, we used one A100 GPU and the HuggingFace Framework with Python 3.10.
@@ 5 Evaluation
5.1 Pre-Training
Perplexity (PPL) (Jelinek et al., 2005) was used as a metric to determine the best performing pre-trained model given the hyper-parameters previously described. Perplexity is defined as the exponentiated average negative log-likelihood of a sequence $X~=~(x_{0},x_{1},x_{2}..x_{t})$ as follows: $PPL(X)=exp\{-\frac{1}{t}\sum^{t}_{i}\ log\ p_{\theta}(x_{i}|x_{<i})\}$. Perplexity was used to evaluate the model’s ability to predict masked tokens on an unseen test dataset representative of the overall corpus distribution. A lower perplexity indicates that the model assigns higher probabilities to the correct next words, i.e., it is more confident in its predictions, less confused by the text.
Table 2 reports the perplexity scores achieved by each model. For conciseness, we only report the best-performing configurations for RooseBERT. Specifically, for both RooseBERT-cont and RooseBERT-scr, a learning rate of $1e^{-4}$ consistently yielded the lowest perplexity on an held out portion of the testing dataset (see Section 3). Final evaluation loss also confirmed this trend. For the cased model, the lowest loss was 0.899 with $1e^{-4}$, followed by 0.901 with $3e^{-4}$, while all other rates (i.e., $5e^{-4}$, $\{1,3,5\}\times e^{-5}$) resulted in losses exceeding 0.925. For the uncased model, the lowest loss was 0.924 with $1e^{-4}$, increasing to 0.933 with $3e^{-4}$, and surpassing 0.951 for the remaining learning rates.
@@ 5.1 Pre-Training
In terms of training time, training the cont models required 18 hours for 150k steps. Training the scr model as considerably more costly, with a total training time of 31 hours and 250k steps.
5.2 Fine-Tuning
BERT, RooseBERT, ConfliBERT and PoliBERTweet available versions (cont, scr, cased and uncased) were fine-tuned on four downstream tasks using the hyper-parameters declared in Section 4. In particular, ConfliBERT is available in both cased and uncased version, with cont and scr configurations. PoliBERTweet is only available in cased version with the cont configuration. The results of the fine-tuning are reported in Table 4. Each reported score represents the average performance over 5 independent runs, each initialised with different seeds (42, 713, 168, 14 and 113).
@@ 5.2 Fine-Tuning
Accuracy
Stance Detection
Macro F1 Score
@@
Macro F1 Score
BERT-base-cased 0.620 $\pm$ 0.007 0.408 $\pm$ 0.003 0.464 $\pm$ 0.003 0.638 $\pm$ 0.025
BERT-base-uncased 0.623 $\pm$ 0.008 0.411 $\pm$ 0.011 0.466 $\pm$ 0.003 0.662 $\pm$ 0.007
ConfliBERT-cont-cased 0.658 $\pm$ 0.004 0.419 $\pm$ 0.008 0.460 $\pm$ 0.002 0.659 $\pm$ 0.006
ConfliBERT-cont-uncased 0.645 $\pm$ 0.006 0.414 $\pm$ 0.005 0.464 $\pm$ 0.003 0.659 $\pm$ 0.007
ConfliBERT-scr-cased 0.644 $\pm$ 0.012 0.422 $\pm$ 0.009 0.459 $\pm$ 0.005 0.621 $\pm$ 0.051
ConfliBERT-scr-uncased 0.633 $\pm$ 0.050 0.415 $\pm$ 0.009 0.458 $\pm$ 0.002 0.659 $\pm$ 0.007
PoliBERTweet 0.596 $\pm$ 0.013 0.407 $\pm$ 0.067 0.443 $\pm$ 0,004 0.667 $\pm$ 0.004
RooseBERT-cont-cased 0.693 $\pm$ 0.007 0.419 $\pm$ 0.012 0.473 $\pm$ 0.001 0.663 $\pm$ 0.007
RooseBERT-cont-uncased 0.699 $\pm$ 0.007 0.425 $\pm$ 0.016 0.466 $\pm$ 0.006 0.611 $\pm$ 0.109
RooseBERT-scr-cased 0.696 $\pm$ 0.004 0.339 $\pm$ 0.021 0.477 $\pm$ 0.005 0.669 $\pm$ 0.011
RooseBERT-scr-uncased 0.702 $\pm$ 0.004 0.424 $\pm$ 0.012 0.480 $\pm$ 0.001 0.674 $\pm$ 0.003
Table 4: Average and standard deviation results of the downstream tasks.
As can be seen, RooseBERT, especially in its scr version, outperforms its competitors across all tasks. More specifically, in sentiment analysis, RooseBERT-scr-uncased achieves a 5% improvement over the best-performing competitor. This gain is statistically significant, as demonstrated by paired t-tests conducted against BERT-base-uncased (p<0.0001) and ConfliBERT-cont-cased (p<0.0001). In the Stance Detection task, RooseBERT-cont-uncased achieves performance comparable to state-of-the-art models, which suggests that, although the VAST dataset is not specifically centered on political debates but on debates in general, RooseBERT can be successfully applied in this context.
In the Argument Component Detection task, RooseBERT-scr-uncased achieves a 2% improvement over best-performing baselines (BERT-base-uncased, p<0.001) and competitors (ConfliBERT-cont-uncased, p<0.001). In the Argument Relation Prediction task, RooseBERT-scr-uncased improves over of 1% over best performing baselines (BERT-base-uncased, p<0.01) and competitors (PoliBERTweet, p<0.05). These results confirm a positive trend induced by the adoption of RooseBERT even for complex natural language understanding tasks such as argument mining.
Overall, the encouraging results obtained on the different tasks emphasise the importance of domain-specific pre-training, where task-relevant embeddings contribute to a more refined understanding of argumentative structure.
5.3 Ablation Study
@@
To better understand which portions of the dataset contribute most to our results, we conducted an ablation study. However, since the dataset consists of eight subsets, performing a leave-one-out ablation would have been computationally expensive. Instead, we clustered the dataset subsets based on the similarity of their content. This was achieved using cosine similarity computed over the average of the embeddings of the sentences of each dataset. To obtain the sentence embeddings we used all-mpnet-base-v2333huggingface.co/sentence-transformers/all-mpnet-base-v2 Sentence Transformer. Figure 1 shows the results of the clustering, performed using Agglomerative Clustering with a distance threshold of 0.2.
Figure 1: Cosine similarity of the pre-training datasets.
@@
As shown in Figure 1, four distinct clusters emerge. One cluster contains only U.S. televised debates (cluster#0), which is expected since they are the only ones not held in a parliamentary setting. Another cluster groups the General Debate of the United Nations (UNGDC) with the United Nations Security Council debates (cluster#1), reflecting their shared institutional origin and likely overlap in terminology and thematic content. The remaining two clusters consist, respectively, of UK Parliament debates on one side (cluster#2), and debates from the European, Australian, Scottish, and Irish legislative chambers (cluster#3).
Once the clusters were identified, we pre-trained four separate RooseBERT models, each using a leave-one-out strategy with respect to the clusters. For instance, the model labelled Cluster#1 was trained on all debates except those in Cluster 1 (i.e., UNGDC and UN Security Council debates). This setup allows us to assess the individual contribution of each cluster to the overall performance by observing the impact of its exclusion.
@@
As shown in Table 5, the exclusion of US Debates appears to have little impact on perplexity, likely because these debates differ in nature from the rest of the dataset, which is predominantly composed of parliamentary debates. On the other side, the exclusion of the European and Australian Parliamentary Debates has the biggest but moderate impact on the perplexity. Yet, all cluster perform worse or as the original model, confirming that most subsets contribute positively to the model’s overall knowledge.
Table 6 shows the results of the fine-tuning of these models on the four different downstream tasks. Results are averaged on a total of 5 runs with different seeds.
Models
Sentiment Analysis
Accuracy
Stance Detection
Macro F1 Score
Arg. Component Detection
Macro F1 Score
Arg. Relation Prediction
Macro F1 Score
Cluster#0 0.706 $\pm$ 0.011 0.415 $\pm$ 0.008 0.475 $\pm$ 0.003 0.667 $\pm$ 0.005
Cluster#1 0.707 $\pm$ 0.005 0.421 $\pm$ 0.009 0.477 $\pm$ 0.003 0.676 $\pm$ 0.005
Cluster#2 0.695 $\pm$ 0.007 0.343 $\pm$ 0.115 0.477 $\pm$ 0.002 0.669 $\pm$ 0.009
Cluster#3 0.706 $\pm$ 0.005 0.411 $\pm$ 0.005 0.477 $\pm$ 0.002 0.663 $\pm$ 0.004
RooseBERT-scr-uncased 0.702 $\pm$ 0.004 0.424 $\pm$ 0.012 0.480 $\pm$ 0.001 0.674 $\pm$ 0.003
Models
Sentiment Analysis
Accuracy
Stance Detection
Macro F1 Score
Arg. Component Detection
Macro F1 Score
Arg. Relation Prediction
Macro F1 Score
gemma-3-4b-it (zero-shot) 0.485 0.232 0.219 0.392
Mistral-7B-Instruct-v0.3 (zero-shot) 0.503 0.302 0.200 0.446
Llama-3.1-8B-Instruct (zero-shot) 0.467 0.395 0.202 0.186
gpt-4.1-nano (zero-shot) 0.488 0.401 0.215 0.427
gemma-3-4b-it (few-shot) 0.484 0.229 0.212 0.354
Mistral-7B-Instruct-v0.3 (few-shot) 0.485 0.302 0.199 0.364
Llama-3.1-8B-Instruct (few-shot) 0.486 0.340 0.197 0.311
gpt-4.1-nano (few-shot) 0.498 0.404 0.213 0.389
RooseBERT’s best model 0.702 0.425 0.480 0.674
Table 7: Performance of LLMs on the downstream tasks in the zero-shot and 3-shot settings.
@@
The performance of the models trained with different dataset configurations remains largely consistent across all downstream tasks, showing minimal to no degradation compared to the original RooseBERT model. This consistency highlights RooseBERT’s strong generalisation capabilities. Notably, we hypothesised a performance drop in the argument component identification and detection, argument relation detection and classification and the sentiment analysis tasks for Cluster#0 and Cluster#2, as these tasks are based on US and UK political debates that were withheld during the training of these models. However, the results show no significant decrease, suggesting that the structural patterns required for these tasks can be effectively learned from other parliamentary debates.
The reader might argue that the results of RooseBERT on downstream tasks are biased by data in the test datasets that is similar to, or identical to, that in the extensive corpora used for pre-training the PLM. However, as shown in Table 6, the results of the ablation study show the opposite, demonstrating that the model can generalise the discourse structure of political debates even when a key domain is withheld during pre-training.
5.4 Experiments with LLMs
Large instruction-tuned language models (LLMs) such as LLaMA-3 (Grattafiori et al., 2024), GPT4 (OpenAI et al., 2024), Mistral 7B (Jiang et al., 2023) or Gemma 3 (Team et al., 2025) have recently been explored for political discourse analysis through prompting strategies and few-shot learning (Törnberg (2023), Kuila and Sarkar (2024)). While full fine-tuning of these models is often computationally prohibitive and unsuitable for resource-constrained environments, zero-shot and few-shot prompting make them a more practical choice for tasks such as sentiment analysis and argument component detection and classification. However, even with prompting, LLMs struggle to capture domain-specific features of political debates, such as implicit argumentation and turn-taking structures.
Table 7 reports the results for gemma-3-4b-it, Mistral-7B-Instruct-v0.3, Llama-3.1-8B-Instruct, and gpt-4.1-nano in the downstream tasks RooseBERT was tested for. All models were tested with a temperature of 0.0 to ensure deterministic outputs. In the Argument Component Detection and Classification task, the models were asked to generate the target sentence with <claim></claim> and <premise></premise> tags while in the other tasks, the models were asked to predict the labels. As can be seen, RooseBERT outperforms LLMs on the four downstream tasks. More specifically, in the stance detection task, the performance of the LLMs is comparable to that of RooseBERT: the VAST dataset used for this task comprises a range of general debatable topics and is not focused on politics, therefore the general-purpose nature of LLMs may be more suitable. However, we observe a performance drop for LLMs compared to RooseBERT when the tasks become more specialized, such as sentiment analysis and argument mining, both in zero-short and few shot settings. Notable errors across tasks included: predicting tags or relations that were not part of the schema (e.g., generating neutral as a sentiment label), hallucinations (such as producing random letters or irrelevant sentences) and disregarding the structure or constraints specified in the prompt. We also notice that the models performed best with sentences comprised between 100 and 400 characters.